In the short run, a competitive firm has a production function,


Q = f(L) = 2.6667L0.75. The output price is $8 per unit and the wage is $5 per hour. Find the short-


run labor demand curve of the firm

1
Expert's answer
2021-11-03T10:41:12-0400

Q = f(L) = 2.6667L0.75

MRPL = MR "\times" MP= w

MR = P = 8

W = 5

MPL = "\frac{\partial Q} {\partial L}"= 2.000025L-0.25

MRPL = 8 "\times" 2.000025L-0.25 = 16.0002L-0.25

16.0002L-0.25 = w

16.0002L-0.25 = 5


"\frac{16.0002}{L^{0.25} }" = 5


16.0002 = 5L0.25


L = 104.86


The short-run labor demand curve of the firm = 104.86